# Arrow-Debreu Meets Kyle: Price Discovery Across Derivatives

### v38

@@ 7.3 Information Efficiency of Prices
Empirically, multi-market price discovery is typically studied through reduced-form decompositions of cointegrated price series with a common efficient component. [^26] See hasbrouck1995one and Gonzalo and Granger (1995). These approaches deliver a relative “who leads” attribution across arbitrage-linked markets rather than an absolute, model-based measure of overall informativeness. Derivative applications require first collapsing option prices to a common implied series. [^27] See, for example, Chakravarty et al. (2004).
@@
Proposition 7.7 characterizes the belief-updating step in the information flow (36). Conditional on aggregate order flow $\omega$, equilibrium prices embed the posterior weight vector $q^{(k)}=(q^{(k)}_{1},\ldots,q^{(k)}_{K})$ (41) over the information types. The logistic-normal form is especially transparent—it is a softmax transformation of a Gaussian “score” vector $Z$. Intuitively, $Z_{\ell}$ aggregates the components of order flow that “look like” signal $s_{\ell}$, while the covariance $(\alpha^{*})^{2}{\bf Q}$ governs which signals are easy to tell apart from order flow. The realized signal $s_{k}$ receives an additional log-odds boost $(\alpha^{*})^{2}$, reflecting that informed order flow systematically tilts $\omega$ toward the true information type.
@@
Kyle’s variance-ratio measure. In Kyle (1985), private information is one-dimensional (a single asset’s value), so a variance ratio for that asset price naturally quantifies how much uncertainty is resolved by trading. In our setting, private information is inherently multi-dimensional: it concerns which payoff distribution (e.g., a higher-moment or tail-risk tilt) governs state prices, and it is learned from joint order flow across many claims. Security-by-security variance ratios are therefore neither aggregative nor invariant—they depend on the particular set of traded claims used to span payoffs and do not capture the cross-claim information spillovers. The posterior-weight measure (43) directly answers the level question posed here—how much weight the equilibrium pricing kernel places on the true payoff distribution.
@@ 8 Empirical Implications
A large empirical literature shows that option prices and order flow embed incremental information about the distribution of the underlying payoff and, in many settings, contribute materially to price discovery. This evidence spans both return-directional information and higher-moment information. [^29] For instance, option-implied variance forecasts subsequent realized variance, and the variance risk premium predicts expected returns (see Bakshi et al. (2003); Carr and Wu (2009); Bollerslev et al. (2009)); option-implied skewness has also been shown to predict cross-sectional returns (see Conrad et al. (2013b)). But in the absence of a formal equilibrium account of how information transmits across option contracts and strategies, empirical work has necessarily been somewhat ad hoc—different studies rely on different predictors and proxies and adopt different contract-specific sorting characteristics. [^30] Examples include Hasbrouck (1995), which considers different price-discovery measures; Bollen and Whaley (2004), which uses net buying pressure to study movements in the implied-volatility surface; Muravyev et al. (2016), which uses option order flow to measure informativeness; and Kaeck et al. (2022), which studies cross-contract impact in index options.
@@
This connects to a large empirical literature that recovers volatility and higher moments from option prices, either through parametric mappings (implied volatility) or through model-free constructions of risk-neutral moments. [^33] See Britten-Jones and Neuberger (2000); Jiang and Tian (2005); Dennis and Mayhew (2002); Bakshi and Madan (2000); Bollen and Whaley (2004); Xing et al. (2010); Cremers and Weinbaum (2010); Bollerslev and Todorov (2011); Bollerslev et al. (2015). While these reduced-form approaches can recover the risk-neutral moments embedded in option prices, inferring future physical moments requires equilibrium identifying restrictions.
@@
The volatility case is a concrete illustration of a broader implication for higher-moment predictability. Combining the informed-demand characterization (37) with the cross price-impact formula in Corollary 7.5, the theory delivers a portfolio-specific forecasting restriction: when informed traders trade a moment-spanning option portfolio, the relevant adverse-selection statistic is the cross price impact between the portfolio’s legs. In Example 7.3, for instance, higher cross price impact between the long–short option positions used to trade right-skewness predicts an increase in the skewness of the underlying return. This logic is consistent with evidence that option-implied skewness can be informative about return asymmetries (e.g., Conrad et al. (2013a)), but existing empirical work typically proceeds through reduced-form moment proxies rather than systematic prediction recipes about where information enters and how it propagates across option markets. We have the following general empirical hypothesis:
@@
Cross price impact between the option legs prescribed by the informed-demand formula (37) predicts subsequent realizations of the corresponding targeted higher moment(s) of the underlying return. [^35] Hypothesis 8.2 can be tested analogously. After estimating $\widehat{\Lambda}$ from the multivariate impact system (45), form the moment-targeting option portfolio prescribed by the informed-demand formula (37)(with legs indexed by $\mathcal{L}^{(m)}$) and compute the within-portfolio cross-impact statistic, $\widehat{CI}^{(m)}_{t}=\!\!\sum_{\begin{subarray}{c}a,b\in\mathcal{L}^{(m)}\\ a\neq b\end{subarray}}\!w^{(m)}_{a}\Big(\widehat{\Lambda}_{a,b}+\widehat{\Lambda}_{b,a}\Big)w^{(m)}_{b}$, where $w^{(m)}$ are the theory-implied leg weights. The hypothesis can then by tested via a predictive regression of the corresponding realized moment over horizon $h$ on the lagged cross-impact statistic, $RM^{(m)}_{t,t+h}=\alpha+\beta\,\widehat{CI}^{(m)}_{t}+\Gamma^{\prime}X_{t}+\varepsilon_{t+h},$ with $\beta$ signed according to whether the prescribed portfolio is long or short the targeted moment.
@@
Systematic factors for the cross-section of option returns can be constructed from the prescribed option portfolios that target higher moments and earn positive returns. [^36] Conjecture 8.4 can be taken to data as follows. First estimate $\widehat{\Lambda}$ from the multivariate impact system (45). Next, for each date $t$ (and within fixed maturity/moneyness buckets), form the theory-prescribed moment-targeting option portfolio from the informed-demand formula (37) for each underlying $i$, with legs indexed by $\mathcal{L}^{(m)}_{i,t}$ and theory-implied weights $w^{(m)}_{i,t}$. Construct the corresponding within-portfolio cross-impact sorting characteristic $\widehat{CI}^{(m)}_{i,t}\equiv\!\!\sum_{\begin{subarray}{c}a,b\in\mathcal{L}^{(m)}_{i,t}\\ a\neq b\end{subarray}}\!w^{(m)}_{a,i,t}\Big(\widehat{\Lambda}_{a,b}+\widehat{\Lambda}_{b,a}\Big)w^{(m)}_{b,i,t}$, and compute the subsequent portfolio return $r^{(m)}_{i,t+1}$. Sort the cross section of prescribed portfolios into quantiles based on $\widehat{CI}^{(m)}_{i,t}$ and form the factor-mimicking portfolio as a high-minus-low spread, $F^{(m)}_{t+1}\equiv\overline{r}^{(m)}_{H,t+1}-\overline{r}^{(m)}_{L,t+1}$, where $H$ and $L$ denote the top and bottom sorting bins. The conjecture predicts $\mathbb{E}[F^{(m)}_{t+1}]>0$ for informational dimensions $m$ along which the prescribed strategies earn positive returns. Finally, the pricing content of $F^{(m)}$ can be evaluated using standard option-asset-pricing tests, e.g., Fama–MacBeth cross-sectional regressions of average returns on estimated factor loadings, with the conjecture implying that exposure to $F^{(m)}$ should help explain cross-sectional variation in expected option returns and reduce pricing errors.
Conjecture 8.4 thus offers candidate option-return factors rooted in information aggregation across contracts. These candidate factors are distinct from (though complementary to) the contract- or stock-level characteristics consider in the existing literature.
@@
(b) displays this insider-induced volatility smiles for different values of true volatility: $0.25$ (solid line, same line as (a)), $0.3$ (dashed line), and $0.35$ (dotted line).
Computation For an order flow realization $\omega$, we use the pricing kernel $P^{*}(,\cdot,|\omega;s_{1})$ to compute the equilibrium option prices across strike $K$. Based on these option prices, we compute the implied volatility $\sigma_{\scriptscriptstyle IV}(\omega,K)$ for each strike $K$. This is repeated for $1,000$ realizations of $\omega$, with the plots displaying the average implied volatilities over log-moneyness. The risk-free rate and dividend yield are set to zero.
@@
A further implication is a sharp comparative static for smile dynamics. As true volatility rises, the implied-volatility curve becomes more convex (see Figure 7b). [^37] As with our other empirical hypotheses, this prediction can be taken directly to data. Fix a maturity $T$. For each day $t$, construct a measure of smile curvature $\kappa_{t}(T)$, e.g., $\kappa_{t}(T)=\tfrac{1}{2}\big(\sigma^{IV}_{t}(-\Delta,T)+\sigma^{IV}_{t}(+\Delta,T)\big)-\sigma^{IV}_{t}(0,T)$ for a fixed delta $\Delta$ (or the quadratic curvature coefficient from a regression of implied volatility on log-moneyness). Estimate $\widehat{\Lambda}_{t}(T)$ from (45) using intraday option midquote changes and signed order imbalance, and form a cross-strike transmission statistic $\widehat{CI}^{\text{wing}}_{t}(T)$ (e.g., an average of symmetrized cross impacts between the two wings, or between each wing and ATM). The model predicts that curvature comoves positively with $\widehat{CI}^{\text{wing}}_{t}(T)$ and steepens in high-volatility states, which can be tested in regressions such as $\kappa_{t}(T)=\alpha+\beta\,\widehat{CI}^{\text{wing}}_{t}(T)+\delta\,\text{Vol}_{t}+\eta\big(\widehat{CI}^{\text{wing}}_{t}(T)\times\text{VolState}_{t}\big)+\Gamma^{\prime}X_{t}+\varepsilon_{t},$ where $\text{Vol}_{t}$ is realized volatility and $X_{t}$ includes standard liquidity controls (spreads, depth, volume, etc.). The intuition is immediate: higher volatility increases the covariance between payoffs at deep ITM and deep OTM strikes, which strengthens cross-strike price impact via the propagation of informed demand across strikes. In this sense, changes in smile curvature are the observable footprint of information aggregation across option markets—rather than a reduced-form feature to be imposed ex ante.
@@ A.1 Model Assumptions
A.5 Proof of Theorem 6.2
@@
This gives the market maker’s posterior (26) in the canonical game.
@@
Suppose $\alpha^{\prime}>0$ solves the equilibrium equation $\Phi(\alpha)=0$, and suppose the market maker holds belief $\alpha^{\prime}{\bf Q}$ in the canonical game. Then, for the insider’s problem (27) conditional on each $s_{k}$ in the canonical game, the first-order condition is sufficient for optimality. Therefore, since the insider’s strategy $\alpha^{\prime}{\bf Q}$ satisfies the collated first-order conditions $\Phi(\alpha^{\prime}){\bf Q}=0$, $\alpha^{\prime}{\bf Q}$ is an equilibrium of the canonical game as defined in (28).
@@
Given the market maker’s belief $\widehat{D}=\alpha^{\prime}{\bf Q}$ in the canonical game, the Hessian matrix $H$ of the insider’s objective function (27) conditional on a signal $s_{k}$ can be computed by differentiating directly his first-order condition (30), which gives
@@
denote the random vector (26) of the market maker’s (normalized) posterior in the canonical game conditional on the insider observing signal $s_{k}$ (with the dependence on $\hat{\omega}=\hat{d}+\widehat{N}$, and $\widehat{D}$ understood). $\bf p$} is a random probability measure on the signal space $S=\{s_{1},\cdots,s_{K}\}$. Let $\alpha>0$ and $Z\stackrel{{\scriptstyle d}}{{\sim}}\mathcal{N}(0,\alpha^{2}{\bf Q})$ be a random vector that is multivariate normal with mean 0 and covariance matrix $\alpha^{2}{\bf Q}$.
@@
Substituting the equilibrium postulate $\widehat{D}=\alpha{\bf Q}$ of (31) into the insider’s first-order condition (30) for the canonical game conditional on (say) $s_{1}$ gives
@@
$\Phi(\alpha)=\mathbb{E}[1-p_{1}-\alpha^{2}p_{1}+\alpha p_{1}^{2}]$
@@
side-by-side gives the matrix equation $\Phi(\alpha){\bf Q}=0$ of (33). By Lemma A.1, an equilibrium is given by a solution $\alpha^{*}>0$ to the equilibrium equation $\Phi(\alpha)=0$.

### v39

@@ 7.3 Information Efficiency of Prices
Empirically, multi-market price discovery is typically studied through reduced-form decompositions of cointegrated price series with a common efficient component. [^26] See Hasbrouck (1995) and Gonzalo and Granger (1995). These approaches deliver a relative “who leads” attribution across arbitrage-linked markets rather than an absolute, model-based measure of overall informativeness. Derivative applications require first collapsing option prices to a common implied series. [^27] See, for example, Chakravarty et al. (2004).
@@
Proposition 7.7 characterizes the belief-updating step in the information flow (29). Conditional on aggregate order flow $\omega$, equilibrium prices embed the posterior weight vector $q^{(k)}=(q^{(k)}_{1},\ldots,q^{(k)}_{K})$ (34) over the information types. The logistic-normal form is especially transparent—it is a softmax transformation of a Gaussian evidence score vector $Z$. Intuitively, $Z_{\ell}$ aggregates the components of order flow that “look like” signal $s_{\ell}$, while the covariance $(\alpha^{*})^{2}{\bf Q}$ governs which signals are easy to tell apart from order flow. The realized signal $s_{k}$ receives an additional log-odds boost $(\alpha^{*})^{2}$, reflecting that informed order flow systematically tilts $\omega$ toward the true information type.
@@
Kyle’s Variance-Ratio Measure
In Kyle (1985), private information is one-dimensional (a single asset’s value), so the prior-posterior variance ratio of that asset price naturally quantifies how much uncertainty is resolved by trading. In our setting, private information is inherently multi-dimensional: it concerns which payoff distribution (e.g., a higher-moment or tail-risk tilt) governs state prices, and it is learned from joint order flow across many claims. Security-by-security variance ratios are therefore neither aggregative nor invariant—they depend on the particular set of traded claims used to span payoffs and do not capture the cross-claim information spillovers. The posterior-weight measure (36) directly answers the level question posed here—how much weight the equilibrium pricing kernel places on the true payoff distribution.
@@ 8 Empirical Implications
A large empirical literature shows that option prices and order flow embed incremental information about the distribution of the underlying payoff and, in many settings, contribute materially to price discovery. This evidence spans both return-directional information and higher-moment information. [^29] For instance, option-implied variance forecasts subsequent realized variance, and the variance risk premium predicts expected returns (see Bakshi et al. (2003); Carr and Wu (2009); Bollerslev et al. (2009)); option-implied skewness has also been shown to predict cross-sectional returns (see Conrad et al. (2013)). But in the absence of a formal equilibrium account of how information transmits across option contracts and strategies, empirical work has necessarily been somewhat ad hoc—different studies rely on different predictors and proxies and adopt different contract-specific sorting characteristics. [^30] Examples include Hasbrouck (1995), which considers different price-discovery measures; Bollen and Whaley (2004), which uses net buying pressure to study movements in the implied-volatility surface; Muravyev et al. (2016), which uses option order flow to measure informativeness; and Kaeck et al. (2022), which studies cross-contract impact in index options.
@@
This connects to a large empirical literature that recovers volatility and higher moments from option prices, either through parametric mappings (implied volatility) or through model-free constructions of risk-neutral moments. [^33] See Britten-Jones and Neuberger (2000); Jiang and Tian (2005); Dennis and Mayhew (2002); Bakshi and Madan (2000); Bollen and Whaley (2004); Xing et al. (2010); Cremers and Weinbaum (2010); Bollerslev and Todorov (2011); Bollerslev et al. (2015). While these reduced-form approaches can recover the risk-neutral moments embedded in option prices, inferring physical moments requires equilibrium identifying restrictions.
@@
The volatility case is a concrete illustration of a broader implication for higher-moment predictability. Combining the informed-demand characterization (30) with the cross price-impact formula in Corollary 7.5, the theory delivers a portfolio-specific forecasting restriction: when informed traders trade a moment-spanning option portfolio, the relevant adverse-selection statistic is the cross price impact between the portfolio’s legs. In Example 7.3, for instance, higher cross price impact between the long–short option positions used to trade right-skewness predicts an increase in the skewness of the underlying return. This logic is consistent with evidence that option-implied skewness can be informative about return asymmetries (e.g., Conrad et al. (2013)), but existing empirical work typically proceeds through reduced-form moment proxies rather than systematic prediction recipes about where information enters and how it propagates across option markets. We have the following general empirical hypothesis:
@@
Cross price impact between the option legs prescribed by the informed-demand formula (30) predicts the corresponding targeted higher moment(s) of the underlying return. [^35] Hypothesis 8.2 can be tested analogously. After estimating $\widehat{\Lambda}$ from the multivariate impact system (38), form the moment-targeting option portfolio prescribed by the informed-demand formula (30)(with legs indexed by $\mathcal{L}^{(m)}$) and compute the within-portfolio cross-impact statistic, $\widehat{CI}^{(m)}_{t}=\!\!\sum_{\begin{subarray}{c}a,b\in\mathcal{L}^{(m)}\\ a\neq b\end{subarray}}\!w^{(m)}_{a}\Big(\widehat{\Lambda}_{a,b}+\widehat{\Lambda}_{b,a}\Big)w^{(m)}_{b}$, where $w^{(m)}$ are the theory-implied leg weights. The hypothesis can then be tested via a predictive regression of the corresponding realized moment over horizon $h$ on the lagged cross-impact statistic, $RM^{(m)}_{t,t+h}=\alpha+\beta\,\widehat{CI}^{(m)}_{t}+\Gamma^{\prime}X_{t}+\varepsilon_{t+h},$ with $\beta$ signed according to whether the prescribed portfolio is long or short the targeted moment.
@@
Systematic factors for the cross-section of option returns can be constructed from the prescribed option portfolios that target higher moments and earn positive returns. [^36] Conjecture 8.4 can be taken to the data as follows. First estimate $\widehat{\Lambda}$ from the multivariate impact system (38). Next, for each date $t$ (and within fixed maturity/moneyness buckets), form the theory-prescribed moment-targeting option portfolio from the informed-demand formula (30) for each underlying $i$, with legs indexed by $\mathcal{L}^{(m)}_{i,t}$ and theory-implied weights $w^{(m)}_{i,t}$. Construct the corresponding within-portfolio cross-impact sorting characteristic $\widehat{CI}^{(m)}_{i,t}\equiv\!\!\sum_{\begin{subarray}{c}a,b\in\mathcal{L}^{(m)}_{i,t}\\ a\neq b\end{subarray}}\!w^{(m)}_{a,i,t}\Big(\widehat{\Lambda}_{a,b}+\widehat{\Lambda}_{b,a}\Big)w^{(m)}_{b,i,t}$, and compute the subsequent portfolio return $r^{(m)}_{i,t+1}$. Sort the cross section of prescribed portfolios into quantiles based on $\widehat{CI}^{(m)}_{i,t}$ and form the factor-mimicking portfolio as a high-minus-low spread, $F^{(m)}_{t+1}\equiv\overline{r}^{(m)}_{H,t+1}-\overline{r}^{(m)}_{L,t+1}$, where $H$ and $L$ denote the top and bottom sorting bins. The conjecture predicts $\mathbb{E}[F^{(m)}_{t+1}]>0$ for informational dimensions $m$ along which the prescribed strategies earn positive returns. Finally, the pricing content of $F^{(m)}$ can be evaluated using standard option-asset-pricing tests, e.g., Fama-MacBeth cross-sectional regressions of average returns on estimated factor loadings, with the conjecture implying that exposure to $F^{(m)}$ should help explain cross-sectional variation in expected option returns and reduce pricing errors.
Conjecture 8.4 thus offers candidate option-return factors rooted in information aggregation across contracts. These candidate factors are distinct from (though complementary to) the contract- or stock-level characteristics considered in the existing literature.
@@
(b) displays the insider-induced volatility smiles for different values of true volatility: $0.25$ (solid line, same line as (a)), $0.3$ (dashed line), and $0.35$ (dotted line).
Computation For an order flow realization $\omega$, we use the pricing kernel $P^{*}(\,\cdot\,|\omega;s_{1})$ to compute the equilibrium option prices across strike $K$. Based on these option prices, we compute the implied volatility $\sigma_{\scriptscriptstyle IV}(\omega,K)$ for each strike $K$. This is repeated for $1,000$ realizations of $\omega$, with the plots displaying the average implied volatilities over log-moneyness. The risk-free rate and dividend yield are set to zero.
@@
A further implication is a sharp comparative static for smile dynamics. As true volatility rises, the implied-volatility curve becomes more convex (see Figure 7b). [^37] As with our other empirical hypotheses, this prediction can be taken directly to the data. Fix a maturity $T$. For each day $t$, construct a measure of smile curvature $\kappa_{t}(T)$, e.g., $\kappa_{t}(T)=\tfrac{1}{2}\big(\sigma^{IV}_{t}(-\Delta,T)+\sigma^{IV}_{t}(+\Delta,T)\big)-\sigma^{IV}_{t}(0,T)$ for a fixed delta $\Delta$ (or the quadratic curvature coefficient from a regression of implied volatility on log-moneyness). Estimate $\widehat{\Lambda}_{t}(T)$ from (38) using intraday option midquote changes and signed order imbalance, and form a cross-strike transmission statistic $\widehat{CI}^{\text{wing}}_{t}(T)$ (e.g., an average of symmetrized cross impacts between the two wings, or between each wing and ATM). The model predicts that curvature comoves positively with $\widehat{CI}^{\text{wing}}_{t}(T)$ and steepens in high-volatility states, which can be tested in regressions such as $\kappa_{t}(T)=\alpha+\beta\,\widehat{CI}^{\text{wing}}_{t}(T)+\delta\,\text{Vol}_{t}+\eta\big(\widehat{CI}^{\text{wing}}_{t}(T)\times\text{VolState}_{t}\big)+\Gamma^{\prime}X_{t}+\varepsilon_{t},$ where $\text{Vol}_{t}$ is realized volatility and $X_{t}$ includes standard liquidity controls (spreads, depth, volume, etc.). The intuition is immediate: higher volatility increases the covariance between payoffs at deep ITM and deep OTM strikes, which strengthens cross-strike price impact via the propagation of informed demand across strikes. In this sense, changes in smile curvature are the observable footprint of information aggregation across option markets—rather than a reduced-form feature to be imposed ex ante.
@@ A.1 Model Assumptions
A.5 Theorem 6.2
A.5.1 Formal Statement of Theorem
The Bayesian trading game between the insider and market maker is isomorphic to the canonical game, defined as follows:
  • There are markets for information contracts $k=1,\cdots,K$.
  • Conditional on signal $s_{k}$, contract $k$ has payoff of $1$ while all others have payoffs of 0.
  • The insider submits orders $\hat{d}\in\mathbb{R}^{K}$ for the information contracts, and the market maker receives order flow $\hat{\omega}=\hat{d}+\widehat{N}$, where $\widehat{N}_{k}$ are i.i.d. $\mathcal{N}(0,1)$ noise trades for $k=1,\cdots,K$.
  • The market maker has a belief $\widehat{D}$ as specified in (22). Upon receiving $\hat{\omega}$, he updates the posterior probability of each signal $s_{k}$ (which is also the contract $k$ price) to
    $\hat{\pi}_{1}(k|\hat{\omega}\kern 0.44998pt;\widehat{D})\propto e^{(\hat{d}^{(k)})^{T}\hat{\omega}-\frac{1}{2}(\hat{d}^{(k)})^{T}\hat{d}^{(k)}}.$ (A.8)
[…]
A.5.2 Proof of Theorem
@@
This gives the market maker’s posterior (A.8) in the canonical game.
@@
A.5.3 Equilibrium in Canonical Game
In the canonical game, an equilibrium is specified by a $K\times K$ strategy matrix $D^{*}=[\delta^{(1)}\;\cdots\;\delta^{(K)}]$ such that, under the same market maker belief $D^{*}$, each strategy $\delta^{(k)}\in\mathbb{R}^{K}$ solves the insider’s problem (A.9) conditional on $s_{k}$, i.e.,
$\delta^{(k)}\in\operatorname*{argmax}_{\hat{d}\,\in\,\mathbb{R}^{K}}J(\hat{d}|k\kern 0.44998pt;D^{*}),\;\mbox{for each}\;k=1,\cdots,K.$ (A.14)
@@ A.5.3 Equilibrium in Canonical Game
Insider FOC in Canonical Game
Let $p\in[0,1]^{I}$ denote the random vector (A.8) of the market maker’s posterior in the canonical game (with the dependence on $\hat{\omega}$ and $\widehat{D}$ understood) and $\mbox{diag}(p)\in\mathbb{R}^{K\times K}$ denote the matrix with $p$ along the diagonal and zeros elsewhere. Then, conditional on observing $s_{k}$ and given market maker belief $\widehat{D}$, the first-order condition for the insider’s problem (A.9) in the canonical game is
[…]
where $e_{k}$ denotes the $k$-th standard basis vector, and the expectation $\mathbb{E}^{(\hat{d};\widehat{D})}$$\cdot\,$] is taken over the distribution of $\hat{\omega}$ induced by the insider’s choice $\hat{d}$ when the market maker holds belief $\widehat{D}$. The terms in the first-order condition (A.15) are the isomorphic counterparts to those in Theorem 5.1.
@@
Suppose $\alpha^{\prime}>0$ solves the equilibrium equation $\Phi(\alpha)=0$, and suppose the market maker holds belief $\alpha^{\prime}{\bf Q}$ in the canonical game. Then, for the insider’s problem (A.9) conditional on each $s_{k}$ in the canonical game, the first-order condition is sufficient for optimality. Therefore, since the insider’s strategy $\alpha^{\prime}{\bf Q}$ satisfies the collated first-order conditions $\Phi(\alpha^{\prime}){\bf Q}=0$, $\alpha^{\prime}{\bf Q}$ is an equilibrium of the canonical game as defined in (A.14).
@@
Given the market maker’s belief $\widehat{D}=\alpha^{\prime}{\bf Q}$ in the canonical game, the Hessian matrix $H$ of the insider’s objective function (A.9) conditional on a signal $s_{k}$ can be computed by differentiating directly his first-order condition (A.15), which gives
@@
denote the random vector (A.8) of the market maker’s (normalized) posterior in the canonical game conditional on the insider observing signal $s_{k}$ (with the dependence on $\hat{\omega}=\hat{d}+\widehat{N}$, and $\widehat{D}$ understood). $\bf p$} is a random probability measure on the signal space $S=\{s_{1},\cdots,s_{K}\}$. Let $\alpha>0$ and $Z\stackrel{{\scriptstyle d}}{{\sim}}\mathcal{N}(0,\alpha^{2}{\bf Q})$ be a random vector that is multivariate normal with mean 0 and covariance matrix $\alpha^{2}{\bf Q}$.
@@
Substituting the equilibrium postulate $\widehat{D}=\alpha{\bf Q}$ of (26) into the insider’s first-order condition (A.15) for the canonical game conditional on (say) $s_{1}$ gives
@@
$\Phi(\alpha)=\mathbb{E}[1-p_{1}-\alpha^{2}p_{1}+\alpha p_{1}^{2}]$ (A.23)
@@
side-by-side gives the matrix equation $\Phi(\alpha){\bf Q}=0$. By Lemma A.1, an equilibrium is given by a solution $\alpha^{*}>0$ to the scalar equilibrium equation $\Phi(\alpha)=0$.